OpenStudy (anonymous):

what is f(x) ? if f(x-1)=x^2+3x+5

OpenStudy (anonymous):

I'm not sure, but I think its (x + 1)^2 + 3(x - 1) + 5

OpenStudy (anonymous):

f(x) = f[(x -1) + 1]

OpenStudy (anonymous):

yup that's right @pratu043 :)

OpenStudy (anonymous):

explain @pratu043

OpenStudy (anonymous):

lol look at what i wrote...f(x-1 + 1) = f(x + 0) = f(x) since f(x-1) = x^2 + 3x + 5 f(x - 1 + 1) = (x+1)^2 + 3(x+1) + 5 got the logic?

OpenStudy (anonymous):

1 and 1 will be cancelled then there will f(x) I didn't get.

OpenStudy (anonymous):

hmmm you get that f(x-1 + 1) = f(x) right?

OpenStudy (anonymous):

now let a = x-1 f(a) = a^2 + 3a + 5 agree?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

now like i said f(x) = f(x-1 + 1) = f(a+1) therefore f(a + 1) = (a+ 1)^2 + 3(a+1) + 5

OpenStudy (anonymous):

oops i made a mistake o.O

OpenStudy (anonymous):

substituting a made me confused -_-

OpenStudy (anonymous):

which mistake.......

OpenStudy (anonymous):

I am also confused ...........

OpenStudy (anonymous):

substituting will actually make it complicated..let me start again do you get that f[(x-1) + 1] = f(x)?

OpenStudy (anonymous):

yes now I am clear thank you...........

OpenStudy (anonymous):

really? lol..great